itsmehjay:

Solve using substitution.. y=2x+7 3x -y = -9

snowflake0531:

\[y=2x+7\] \[3x-y=-9\] Can you plug the first equation into the second?

itsmehjay:

Oh yeah uhhh 3x - (2x + 7) = -9??

snowflake0531:

Yep!!!!! so now distribute out the parenthesis, and then solve for x

itsmehjay:

2x + 7 = 9x?

itsmehjay:

3x - 9x = -6x..

snowflake0531:

Where are you getting this lol \[3x-(2x+7)=-9\] Imagine the negative sign as negative 1 And do the distributive property So, the signs will change and become \[3x-2x-7=-9\] now can you solve for x?

itsmehjay:

oh ok so x - 7 = -9 x = -9 + 7 x = -2? o:

snowflake0531:

Yep So now plug in the x into one of the equations to get y

itsmehjay:

I'll use the first one y=2 (-2) + 7

itsmehjay:

y= -4 + 7 y= 3

snowflake0531:

Yep~

itsmehjay:

Wowie your smort thank you so much >.<

snowflake0531:

why, thank you xd and yw~